OpenStudy (anonymous):

use implicit differentiation to find y' if x^3-xy+y^3=7

OpenStudy (anonymous):

3x^2 - y - xy' + 3y^2y' = 0 Now solve for y'. Your answer will have x's and y's in it.

zepdrix (zepdrix):

Mark do you understand how bang came up with those middle two terms? It requires the product rule:\[\Large (-xy)'\quad=\quad -(x)'y-x(y)'\quad=\quad -(1)y-x(y')\]

OpenStudy (anonymous):

Yeah, i think the product rule with implicit differentiation confuses a lot of people.

OpenStudy (anonymous):

how do i get rid of the y' in the exponent

zepdrix (zepdrix):

Hmm there are no y' in exponent positions.\[\Large (y^3)'\quad=\quad (3y^2)y'\]

OpenStudy (anonymous):

I got it now i was reading it wrong before

zepdrix (zepdrix):

Oh ok :)